Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 1,230 hours each month to produce 2,050 sets of covers. The standard costs associated with this level of production are: Total Per Set of Covers Direct materials $ 27,675 $ 13.50 Direct labor $ 8,610 4.20 Variable manufacturing overhead (based on direct labor-hours) $ 4,920 2.40 $ 20.10 During August, the factory worked only 750 direct labor-hours and produced 1,500 sets of covers. The following actual costs were recorded during the month: Total Per Set of Covers Direct materials (3,900 yards) $ 19,500 $ 13.00 Direct labor $ 6,600 4.40 Variable manufacturing overhead $ 4,800 3.20 $ 20.60 At standard, each set of covers should require 1.8 yards of material. All of the materials purchased during the month were used in production. Required: 1. Compute the materials price and quantity variances for August. 2. Compute the labor rate and efficiency variances for August. 3. Compute the variable overhead rate and efficiency variances for August. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

Material Price Variance = (Standard Price – Actual Price)*Actual Quantity Purchased

= (7.5 – Actual Rate)*3,900

= 7.5*3900 – 19,500

= $9,750 Favorable

Material Quantity Variance = (Standard Quantity – Actual Quantity Used)*Standard Price

= (1,500*1.8 – 3,900)*7.5

= $9,000 Unfavorable

Labor Rate Variance = (Standard Rate – Actual Rate)*Actual Hours

= (4.2*1.66666 – Actual Rate)*750

= $1,350 Unfavorable

Efficiency Variance = (standard hours – actual hours)*Standard Rate

= (1500/1.66666 – 750)*7

= $1,050 favorable

Variable overhead rate variance = (Standard Rate – Actual Rate)*Actual Hours

= (2.4*1.6666 – Actual Rate)*750

= $1,800 U

Efficiency Variance = (standard hours – actual hours)*Standard Rate

= (150)*4

= $600 F